1. Briefly explain why the heat capacity of the calorimeter must be calculated in this experiment.

The experiment includes the calibration of the calorimeter, the measurement of temperatures to determine enthalpy change. Chemicals include ammonium hydroxide, hydrochloric acid and ammonium chloride.
Answer #1

The calorimeter calibration equation is given by

Q= C×( change in temperature when a substance is burnt inthe calorimeter)

Where C= calorimeter constant which is nothing but the heat capacity of the calorimeter and can be found usous following equation

C= Q/ ( change in temperature)

In order to find Q we need to know the Q and Q can be found from the equation follows

Q= I×V× T×I= amount of heat electrically given to the calorimeter.

It is clear from the above argument that in order to calibrate the calorimeter we must find the heat capacity first.